When launched, H.M.S. Titanic was the most passive movable object ever built by humans, having a mass of 6.0 × 107kg. What would have been the magnitude of the net force required to give Titanic an acceleration of magnitude 0.2 m s-2?

Grade:12th pass

3 Answers

Rakhi Bisht
7 years ago
Mass of Titanic = 6.0 × 107 kg
Acceleration, α = 0.2 m s-2
Or F = 6.0 × 107 × 0.2 N = 1.2 × 107 N = 12 × 106 N = 12 million newton.
Satendra Kumar
7 years ago
Hi Kumar,
below the answer for your question:
we have mass = 6.0 × 10^7kg, Acceleration, α = 0.2ms^-2 or F = 6.0 × 10^7 × 0.2N = 1.2 × 10^7N = 12 × 10^6N = 12 million Newton.
Siddharth Arora
7 years ago
 
Mass of Titanic = 6.0 × 107 kg
Acceleration, α = 0.2 m s-2
Or F = 6.0 × 107 × 0.2 N = 1.2 × 107 N = 12 × 106 N = 12 million newton.